OpenStudy (anonymous):

Which of the following represents the greatest energy transition from a higher energy level to a lower one? A. emission of a red photon of 770 nm B. emission of a green photon of 550 nm C. emission of a yellow photon of 590 nm D. emission of a blue photon of 450 nm

OpenStudy (michele_laino):

here we have to compute four energies, namely we have to apply this formula: \[E = \frac{{hc}}{\lambda }\]

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

so case red photon: \[E = \frac{{hc}}{\lambda } = \frac{{6.62 \times {{10}^{ - 34}}3 \times {{10}^8}}}{{770 \times {{10}^{ - 9}}}} = ...joules\]

OpenStudy (anonymous):

i am confused... what does this mean? 10^-34 3 ?

OpenStudy (michele_laino):

oops.. sorry: \[E = \frac{{hc}}{\lambda } = \frac{{6.62 \times {{10}^{ - 34}} \times 3 \times {{10}^8}}}{{770 \times {{10}^{ - 9}}}} = ...joules\]

OpenStudy (anonymous):

oh! okie! so we get 2.579E-19

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

does that mean choice A is our soluition?

OpenStudy (michele_laino):

no, we have to to the same computation for other wavelength. case green photon: \[E = \frac{{hc}}{\lambda } = \frac{{6.62 \times {{10}^{ - 34}} \times 3 \times {{10}^8}}}{{550 \times {{10}^{ - 9}}}} = ...joules\]

OpenStudy (anonymous):

ok! so we get 3.61E-19

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

case yellow photon: \[E = \frac{{hc}}{\lambda } = \frac{{6.62 \times {{10}^{ - 34}} \times 3 \times {{10}^8}}}{{590 \times {{10}^{ - 9}}}} = ...joules\]

OpenStudy (anonymous):

and we get 3.366E-19!

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

finally, case blue photon: \[E = \frac{{hc}}{\lambda } = \frac{{6.62 \times {{10}^{ - 34}} \times 3 \times {{10}^8}}}{{450 \times {{10}^{ - 9}}}} = ...joules\]

OpenStudy (anonymous):

4.42E-19!

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

reassuming, we have these subsequent energies: 2.579E-19 J, 3.61E-19 J, 3.366E-19 J, 4.42E-19 J. which is the greatest one?

OpenStudy (anonymous):

the last! so our solution is choice D?

OpenStudy (anonymous):

oh wait oops sorry it is the first, right? so chocie A is our solution?

OpenStudy (michele_laino):

I think that 4.42 is greater than 2.579

OpenStudy (anonymous):

but it is negative?

OpenStudy (anonymous):

E-19? :/

OpenStudy (michele_laino):

that is an exponential, all our energies have 10^(-19) as exponential

OpenStudy (anonymous):

oh so our solution is still choice D?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

yay! thanks:)

OpenStudy (michele_laino):

:)